OpenStudy (mariebrittany07):

Can someone help me please You work for 4 hours on Saturday and 8 on Sunday.You also receive a $50 bonus.You earn $164. How much did you earn per hour?

OpenStudy (kirbykirby):

The bonus is not included in your hourly wage. It is simply an extra that you earn. Thus, we can say 164 - 50 = $114 in hourly wages. If we assume the hourly wage is always the same, then you worked for 8+4 = 12 hours. So, If 12 hours = $114, then 1 hour = ? Well, $114/12 hours = $9.5/hour

OpenStudy (anonymous):

You would make $9.50 cent. For you to earn $164 with a bonus of $50 and having to work for four hours on Saturday and eight on Sunday.

OpenStudy (mariebrittany07):

okay thanks a million